# Benjamin Carleton

### v6

@@ 5 Conclusion
In this paper, for all collapsing electoral control types found in Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [${}^{+}$22]—those are all such collapses for the domains and groupings studied there (see [HHM20] and especially [${}^{+}$22])—we proved that even the search-problem complexities of the collapsing types are polynomially equivalent (given oracle access to the election system’s winner problem).
In doing this—and building on earlier notions of relating search problems—we defined reductions that for the case of collapsing electoral control problems express how solutions to one can be efficiently converted to solutions to the other.
Also, for the key concrete systems plurality, veto, and approval, we establish—as polynomial-time computable or as ${\rm SAT}$-equivalent—the complexity of their (now collapsed) search problems. An interesting open direction would be, for that part of our work, to seek more general results, such as dichotomy theorems covering broad collections of election systems. However, that may be difficult since not much is known as to dichotomy theorems even for the decision cases of (unweighted) control problems (however, see [HHS14, HS16, FHH15]), though the few known such cases would be natural starting points to look at in this regard.

### v7

@@ 5 Conclusion
In this paper, for all collapsing electoral control types found in Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [CCH+24]—those are all such collapses for the domains and groupings studied there (see [HHM20] and especially [CCH+24])—we proved that even the search-problem complexities of the collapsing types are polynomially equivalent (given oracle access to the election system’s winner problem).
In doing this—and building on earlier notions of relating search problems—we defined reductions that, for the case of collapsing electoral control problems, express how solutions to one can be efficiently converted to solutions to the other.
Also, for the key concrete systems plurality, veto, and approval, for each of their collapsing pairs we establish either that both of the pair’s search versions are polynomial-time computable, or that both of the pair’s search versions are $\rm SAT$}-equivalent (see Table 1). An interesting open direction would be, for that part of our work, to seek more general results, such as dichotomy theorems covering broad collections of election systems. However, that may be difficult since not much is known as to dichotomy theorems even for the decision cases of (unweighted) control problems (however, see [HHS14, HS16, FHH15]), though the few known such cases would be natural starting points to look at in this regard.
Acknowledgments
Preliminary versions of this work appeared in Proceedings of the 22nd International Conference on Autonomous Agents and Multiagent Systems [CCH+23a] and in Proceedings of the 21st European Conference on Multi-Agent Systems [CCH+25]. We thank the anonymous referees for their helpful comments. This work is supported in part by NSF grants CCF-2006496 and DUE-2135431, CIFellows grant CIF2020-UR-36, and a Renewed Research Stay grant from the Alexander von Humboldt Foundation.